OpenStudy (anonymous):

f(x)=2x-1and g(x)=sqrt(x), find g(f(x))

OpenStudy (anonymous):

\[g(f(x))=g(2x-1)=\sqrt{2x-1}\]

OpenStudy (anonymous):

@satellite73 why is g(2x-1) sqrt(2x-1) instead of 2*sqrt(x)-1?

OpenStudy (hba):

@livelaughlilz You have to substitute the value of \[f(x)=2x-1\] In place of x in \[g(x)=\sqrt(x)\]

OpenStudy (anonymous):

oh so the "x" becomes "2x-1"

OpenStudy (hba):

Exactly :)

OpenStudy (anonymous):

thanks